# Debias Can be Unreliable: Mitigating Bias Issue in Evaluating Debiasing Recommendation

### v1

@@ 5 URE scheme
Considering the impractical accessibility of $D_{full}$, the pivotal role of $D_{rand}$ emerges in enabling an unbiased estimation of Recall@$K$ on $D_{full}$. Nevertheless, as mentioned above, our contention is that the traditional evaluation scheme on $D_{rand}$ exhibits bias. In response to this challenge, we introduce a pioneering evaluation scheme coined URE, delineated in Figure 4. URE relies solely on $D_{rand}$ to achieve an unbiased estimation of Recall@$K$ on $D_{full}$, which consists of the following steps:
@@
  • Compute the $\widehat{\text{Recall@}K}$ for this user:
    $\widehat{\text{Recall@}K}=\frac{m}{n},$
    where $n$ denotes the number of items with positive label in Sorted $D_{rand}$, and $m$ denotes the number of these items with higher $\hat{y}$ than the ($K$+1)-th item.
@@
$E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall@}K}\right]=\text{Recall}@K,$ (4)
@@
The proof can be found in Appendix A.4. This signifies that the estimates derived from URE are unbiased.
In addition to the theoretical analysis, we provide an intuitive explanation to show how our URE can unbiasedly estimate the Recall@$K$ on $D_{full}$. The core reliance is that $D_{rand}$ represents a random selection from the sorted list of $D_{full}$. Hence, the ratio of items with positive label in $D_{rand}$ which have a higher $\hat{y}$ than the ($K$+1)-th item can be used to unbiasedly estimate the positive ratio of $D_{full}$. Our proposed URE scheme, despite its apparent simplicity, can generate outstanding predictive accuracy. This is due to the models’ prediction information of all candidate items, which is not considered by traditional evaluation scheme. We will demonstrate the effectiveness of this scheme in following sections.
Figure 5: (a) The Kendall rank coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ as well as the conventional Recall@$\overline{K}$. (b) The Kendall rank coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$. We mark $K_{max}$ when every curve reaches its highest coefficient. The rank correlations in Figure (a) and (b) are statistically significant with $p<$ 0.01.
6 Experiments
Table 4: The evaluation results by using the URE scheme. We also show the Recall@$\overline{5}$ under the traditional evaluation scheme for comparison. The best results are in bold and the second best are underlined. Notably, the Recall@$5$ metric is not available in Yahoo!R3, as it lacks a fully-exposed dataset. The results are statistically significant with $p<$ 0.01.
Backbone Model KuaiRec Yahoo!R3
Traditional scheme Fully-Exposed Dataset URE scheme Fully-Exposed Dataset Traditional scheme URE scheme
Recall@$\overline{5}$ Recall@$5$ $\widehat{\text{Recall@5}}$ $\widehat{\text{Recall@500}}$ Recall@500 Recall@$\overline{5}$ $\widehat{\text{Recall@5}}$ $\widehat{\text{Recall@500}}$
MF Bias 0.3546 0.0215 0.0213 0.3280 0.3237 0.7085 0.0131 0.6777
IPS 0.3102 0.0213 0.0230 0.2771 0.2743 0.6572 0.0095 0.6236
DR 0.5465 0.0353 0.0362 0.5077 0.5050 0.7195 0.0075 0.6879
AutoDebias 0.5482 0.0355 0.0352 0.5129 0.5097 0.8085 0.0550 0.7849
NegS 0.5358 0.0327 0.0326 0.5052 0.5012 0.8330 0.0490 0.8001
LightGCN Bias 0.4859 0.0257 0.0267 0.4503 0.4482 0.7883 0.0491 0.7680
IPS 0.4889 0.0266 0.0273 0.4539 0.4522 0.7584 0.0457 0.7367
DR 0.4849 0.0262 0.0267 0.4500 0.4457 0.7576 0.0455 0.7373
AutoDebias 0.5181 0.0317 0.0323 0.4571 0.4558 0.8046 0.0592 0.7820
NegS 0.5288 0.0237 0.0245 0.4954 0.4944 0.8141 0.0616 0.7913
In this section, we conduct experiments to verify the effectiveness of the URE scheme and re-evaluate the debiasing methods using the proposed URE scheme.
6.1 Unbiasedness Analysis
In this subsection, we test the AutoDebias model using the KuaiRec dataset to verify that the proposed $\widehat{\text{Recall@}K}$ in the URE scheme is an unbiased estimation of Recall@$K$ on $D_{full}$.
In Figure 5a, we investigate whether there is order preservation between Recall@$K$ and our $\widehat{\text{Recall@}K}$ as well as the conventional Recall@$\overline{K}$. Concretely, we randomly select model’s hyperparameter combinations to train the AutoDebias model on the Kuairec dataset. Then we calculate the Kendall rank correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$, as well as Recall@$\overline{K}$ to demonstrate the ranking consistency. As shown in Figure 5a, when $K$ is small (e.g. $K=30$), both Recall@$\overline{1}$ and Recall@$\overline{5}$ exhibits poor correlations with Recall@$K$. In contrast, the Kendall correlation coefficient between the estimated $\widehat{\text{Recall@}K}$ and Recall@$K$ is higher than 0.9. This means that the model we choose based on $\widehat{\text{Recall@}K}$ is more likely to perform better on Recall@$K$ on $D_{full}$, which further demonstrates the effectiveness of the URE scheme.
In Figure 5b, we report the Kendall rank correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$. As we desired, only when K is the same, Recall@$K$ and $\widehat{\text{Recall@}K}$ have the highest correlation coefficients. That is, URE allows us to select the best debiased model based on the Recall@$K$ with a desired $K$.
Based on our discussion, we show that our URE scheme unbiasedly estimates Recall@$K$ for different values of $K$. Moreover, the high Kendall coefficients between $\widehat{\text{Recall@}K}$ and Recall@$K$ ensure an accurate debiased model selection.
6.2 Re-evaluation Results
In this subsection, we re-evaluate some classical debiasing methods, and report the results in Table 4. We have the following key findings:
  • Based on the results on KuaiRec dataset, we observe that the evaluation result $\widehat{\text{Recall@5}}$ of the URE scheme aligns closely with the true Recall@5 on the fully-exposed dataset. This observation suggests a strong unbiasedness in the URE scheme. In contrast, the evaluation results of the traditional scheme exhibit a significant discrepancy from the true Recall@5, and it fails to ensure the preservation of order in the results.
  • In some instances, the $\widehat{\text{Recall@5}}$ estimated by the URE scheme is inversely ordered compared to the actual Recall@5. This phenomenon can occur when the actual Recall@5 values are closely aligned, as is the case with the MF-based DR method at 0.353 and the MF-based AutoDebias method at 0.355. We contend that such a degree of deviation is within an acceptable range.
  • When using $\widehat{\text{Recall@5}}$ to verify the debiasing performance, some debiasing methods perform worse than the Bias model, which encourages the development of further debiasing techniques.
  • Based on the proposed URE, the performance of NegS is sub-optimal as expected in most instances, which implies that URE possesses potential merits in selecting truly effective debiasing approaches. In the LightGCN-based method on Yahoo!R3, the Recall@5 of NegS exceeds that of all other methods, which we attribute to the fact that these methods are not successfully debiasing in this case.
  • There is an inconsistency in the ranking results based on $\widehat{\text{Recall@}K}$ metric with large and small $K$ values. This means we should develop different debiasing methods to accommodate Recall@$K$ with different $K$ on $D_{full}$.
7 Related Work
7.1 Debiasing Recommendation
Recent recommendation debiasing strategies can be divided in two main categories: data-level debiasing and representation-level debiasing. At the data level, debiasing is always achieved by inverse propensity reweighting  (Schnabel et al. 2016; Wang et al. 2016; Li et al. 2023c) and doubly robust learning (Dai et al. 2022; Li et al. 2023b; Li, Zheng, and Wu 2023). At the representation level, information bottleneck (Wang et al. 2020), causal representation learning (Liu et al. 2020; Yang et al. 2023), and invariant representation learning (Wang et al. 2022; Yang et al. 2023; Pan et al. 2023) are widely used. For evaluating the debiasing effectiveness of these methods, researchers often evaluate Recall@$\overline{K}$ performance (e.g., Recall@$\overline{5}$) of the learned recommendation models on publicly available randomly-exposed data (Chen et al. 2023) — Coat, Yahoo!R3 and KuaiRand. However, we point out such an evaluation scheme on $D_{rand}$ may be problematic.
7.2 Sampling Metrics
The traditional evaluation scheme that calculates Recall@$\overline{K}$ metric on $D_{rand}$ can be thought of as a type of sampling metric. (Krichene and Rendle 2020) is the first work to highlight the inconsistency between sampled metrics on the sampled data and exact metrics on full data. It also proposes a method based on the sampled data to unbiasedly estimated Recall@$K$ metrics on full data. The simple estimation method suffers from the problem of large variance. Subsequent work (Li et al. 2023a, 2020; Jin et al. 2021; Dallmann, Zoller, and Hotho 2021) improves the estimation scheme and reduce the estimation variance. However, all the works mentioned above only focus on the case where the test set contains only one positive sample. In this way, the conclusions are not applicable to the debiasing task in that the test set contains multiple positive samples. Furthermore, there are no discussions about the effect of traditional evaluation scheme on the debiasing task.
8 Conclusion And Future Work
In this paper, we reviewed traditional evaluation scheme on the randomly-exposed datasets by comparing it with the gold standard evaluation on the fully-exposed datasets. According to the theoretical and empirical analyses, we revealed the inconsistency between them w.r.t. the values of $\overline{K}$ and $K$. We conducted an entrapment study, where a hard negative sampling trick without any debiasing techniques can achieve incredible debiasing performance by using the traditional evaluation scheme on the randomly-exposed datasets, further validating the traditional evaluation scheme on the randomly-exposed datasets may not reflect true debiasing performance and results in unreliable conclusions. Accordingly, we designed the URE scheme to unbiasedly estimate the true Recall performance of the fully-exposed datasets. Both theoretical and empirical analyses have validated the effectiveness of our proposal.
In future work, we will investigate the unbiased estimation of other Top-$K$ metrics, like Precision and NDCG on the fully-exposed dataset to further extend our conclusions. Besides, improving URE from the perspective of bias-variance balancing is another interesting direction. We hope that the insight of this paper could bring new inspiration for debiasing recommendations.
Appendix A Appendix
A.1 Datasets
We use two real-world datasets (Yahoo!R3 and KuaiRec) for our experiments.
  • Yahoo!R3: This dataset is widely-used in debiasing literature which consists of a biased dataset and a randomly-exposed dataset collected by a random logging strategy. We follow (Chen et al. 2021) and randomly sample 10% of the randomly-exposed data for debiasing training (required by AutoDebias and IPS), and take the remaining 90% as $D_{rand}$ for evaluating. The interactions between users and items are ratings ranging from one to five. The ratings larger than three are labeled as positive ($r_{u,i}$ = 1), otherwise negative ($r_{u,i}$ = 0). Notably, the Recall@$K$ metric is not available in Yahoo!R3, as it lacks a fully-exposed dataset. However, in our proposed URE scheme, we can use $D_{rand}$ to compute $\widehat{\text{Recall@}K}$ for unbiased estimation of Recall@$K$.
  • KuaiRec: This dataset is the only existing fully-exposed dataset. We construct three types of datasets from this original fully-exposed data: biased data for training, fully-exposed data $D_{full}$ for evaluating, and randomly-exposed data $D_{rand}$ for evaluating. Concretely, we sample items for each user inversely proportional to their video duration and combine them with their ratings to obtain biased data. Interactions on the $k$-th shortest video are sampled with probability $p=e^{k/\lambda}/\sum_{k}e^{k/\lambda}$, $\lambda$ controls the degree of bias towards short videos. The remaining items and ratings is considered as the fully-exposed data ($D_{full}$) for evaluating. Following (Marlin and Zemel 2009), we construct randomly-exposed data ($D_{rand}$) by uniformly sampling $\overline{N}$ items for each user on $D_{full}$. The watch_ratio is binarized with threshold 2 following (Gao et al. 2022a). That is, the video watching ratio larger than 2 is labeled as positive ($r_{u,i}$ = 1), otherwise negative ($r_{u,i}$ = 0).
A.2 Proof of Theorem 3
@@
$E_{\mathcal{R}}\left[\text{Recall}@K\right]=\sum_{x=1}^{min(K,N^{+})}\frac{x}{N^{+}}\frac{C^{x}_{N^{+}}C_{N-N^{+}}^{K-x}}{C_{N}^{K}}$ (5)
$=\sum_{x=1}^{min(K,N^{+})}\frac{xN^{+}K}{xN^{+}N}\frac{C^{x-1}_{N^{+}-1}C^{(K-1)-(x-1)}_{(N-1)-(N^{+}-1)}}{C_{N-1}^{K-1}}$ (6)
$=\frac{K}{N}\left[\sum_{x=1}^{min(K,N^{+})}\left[C_{N^{+}-1}^{x-1}C_{(N-1)-(N^{+}-1)}^{(K-1)-(x-1)}/C^{K-1}_{N-1}\right]\right]$ (7)
$=\frac{K}{N}\left[\ C_{N-1}^{K-1}/C_{N-1}^{K-1}\right]=\frac{K}{N}.$ (8)
$E_{\mathcal{R}}\left[E_{\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}]\right]=E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right].$ (9)
$E_{{N^{+}}^{\prime}}\left[E_{\mathcal{R},\mathcal{S}_{\overline{N}}}[\text{Recall@}\overline{K}|{N^{+}}^{\prime}]\right]$ (10)
$=E_{{N^{+}}^{\prime}}\left[\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}\frac{x}{{N^{+}}^{\prime}}\frac{C^{x}_{{N^{+}}^{\prime}}C^{\overline{K}-x}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}}^{\overline{K}}}\right]$ (11)
$=E_{{N^{+}}^{\prime}}\left[\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}\frac{x\overline{K}{N^{+}}^{\prime}}{x{N^{+}}^{\prime}\overline{N}}\frac{C^{x-1}_{{N^{+}}^{\prime}-1}C^{(\overline{K}-1)-(x-1)}_{N-{N^{+}}^{\prime}}}{C_{N-1}^{\overline{K}-1}}\right]$ (12)
$=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\frac{\sum_{x=1}^{min(\overline{K},{N^{+}}^{\prime})}C^{x-1}_{{N^{+}}^{\prime}-1}C^{(\overline{K}-1)-(x-1)}_{\overline{N}-{N^{+}}^{\prime}}}{C_{\overline{N}-1}^{\overline{K}-1}}\right]$ (13)
$=E_{{N^{+}}^{\prime}}\left[\frac{\overline{K}}{\overline{N}}\right]=\frac{\overline{K}}{\overline{N}}=\frac{K}{N}=E_{\mathcal{R}}[\text{Recall@}K].$ (14)
A.3 Proof of Theorem 2
Proof.
It has been proven that when optimizing the BPR loss function using uniform negative sampling, the optimization objective of the model is AUC in the previous study (Rendle et al. 2009). Therefore, we only need to prove that the BCE loss function commonly used in debiasing tasks is an upper bound of the BPR loss function. Specifically, the BCE loss function is:
$\delta_{BCE}=\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\left[-\log(\sigma(f_{\theta}(u,i))-\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}\log(1-\sigma(f_{\theta}(u,j))\right],$ (15)
where $\sigma(x)=1/(1+e^{-x})$, $f_{\theta}(u,i)$ denotes a model with trainable parameter $\theta$ for predicting the rating of item $i$ by user $u$, and $\mathcal{D}$ refers to the train dataset containing ($u$, $i$) pairs.
BPR loss function:
$\delta_{BPR}=\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log(\sigma(f_{\theta}(u,i)-f_{\theta}(u,j)),$ (16)
where $\mathcal{I}$ denotes the item set.
By transforming the BCE loss function, we can obtain:
$\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\left[-\log(\sigma(f_{\theta}(u,i))-\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}\log(1-\sigma(f_{\theta}(u,j))\right]$ (17)
$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log(\sigma(f_{\theta}(u,i)\cdot(1-\sigma(f_{\theta}(u,j)))))$ (18)
$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\left[\frac{1}{1+e^{-f_{\theta}(u,i)}}\cdot\frac{e^{-f_{\theta}(u,j)}}{1+e^{-f_{\theta}(u,j)}}\right]$ (19)
$=$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\frac{1}{1+e^{-f_{\theta}(u,i)}+e^{f_{\theta}(u,j)}+e^{-f_{\theta}(u,i)+f_{\theta}(u,j)}}$ (20)
$\geq$ $\sum_{(u,i)\in\mathcal{D}}\frac{1}{|\mathcal{D}|}\frac{1}{|\mathcal{I}|}\sum_{j\in\mathcal{I}}-\log\frac{1}{1+e^{-(f_{\theta}(u,i)-f_{\theta}(u,j))}}=\delta_{BPR}.$ (21)
Thus, we have proved that the commonly used BCE loss function is an upper bound of the BPR loss function. Therefore, optimizing the BCE loss function is equivalent to optimizing the upper bound of the BPR loss function, which equals optimizing AUC.
Based on the aforementioned mentioned equivalence and referring to Theorem 1 and Theorem 3 from (Shi et al. 2023), we derive that optimizing BCE loss enhanced with DNS($H$, $U$) approximates the optimization of the Recall@$K$ metric. Specifically, the larger the $H/U$ is, the Recall@$K$ metric with a larger $K$ on $D_{full}$ is optimized. ∎
A.4 Proof of Theorem 3
Proof.
@@ A.4 Proof of Theorem 3
Since Recall is only related to positive samples, we can only consider positive samples. Let’s assume that in a fully exposed scenario, there are a total of $N$ positive samples, out of which $M$ samples are ranked higher than the sample ranked at $K+1$ in Sorted ItemID in Figure 4. Assuming that we have randomly sampled $\overline{N}$ samples, which have $n$ positive samples and the rest are negatives. Since the sampling process is random, the probability of the number $m$ of positive samples that rank higher than the sample ranked at $K+1$ in Sorted ItemID can be calculated as $(C^{m}_{M}C^{n-m}_{N-M}/C^{n}_{N})$, thus we have:
@@
$E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall}@K}\right]$ $=E_{n^{\prime}}\left[E_{\mathcal{S}_{\overline{N}}}[m/n|n^{\prime}]\right]$ (22)
$=E_{n^{\prime}}\left[\sum_{{m=1}}^{min(M,n^{\prime})}\frac{m}{n^{\prime}}\frac{C^{m}_{M}C^{n^{\prime}-m}_{N-M}}{C^{n^{\prime}}_{N}}\right]$ (23)
$=E_{n^{\prime}}\left[\sum_{{m=1}}^{min(M,n^{\prime})}\frac{mn^{\prime}M}{mn^{\prime}N}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$ (24)
$=E_{n^{\prime}}\left[\frac{M}{N}\sum_{{m=1}}^{min(M,n^{\prime})}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$ (25)
$=M/N=\text{Recall}@K.$ (26)
A.5 Compared Methods
We utilize two representative recommendation models, Matrix Factorization (MF) (Koren, Bell, and Volinsky 2009) and graph neural network model LightGCN (He et al. 2020), as the backbones. Based on the Theorem 2, We devise NegS, which optimizes the backbone model through BCE loss accompanied by the hard negative sampling, and compare NegS with following debiasing methods:
  • IPS (Schnabel et al. 2016) weights the loss of each sample with an inversed propensity. We follow (Chen et al. 2021) and calculate the propensity with naive bayes.
  • DR (Wang et al. 2019) integrates imputed errors and propensities while mitigating the impact of high propensity variance.
  • AutoDebias (Chen et al. 2021) propose a meta-learning-based algorithm to learn debiasing parameters from a small set of randomly-exposed data.
A.6 Metrics
We use Recall metric to evaluate models for our following experiments. The Recall metric is the most important metric in the candidate generation stage of industrial recommendation systems. It is also a commonly used metric in research on debiasing recommendations. In future work, we will investigate the unbiased estimation of other Top-$K$ metrics, like Precision and NDCG on the fully-exposed dataset to further extend our conclusions.
A.7 Setup
All the models are optimized by Adam optimizer. We use grid search to find the best hyperparameters. The search space of learning rate and weight decay are [1e-4, 1e-3, 1e-2, 1e-1]. For the proposed NegS, we tune $H/U$ and $U$ within the ranges of [0.1,0.2,0.3,0.4,0.5,0.6,0.7,0.8,0.9,1] and [50,100,200] respectively. For details on the dataset, see Appendix A.1. For details on the baseline, see Appendix A.5. For details on the matrics, see Appendix A.6.

### v2

@@ 4. URE scheme
In this section, we introduce a pioneering evaluation scheme coined URE, delineated in Figure 2. URE relies solely on $D_{rand}$ to achieve an unbiased estimation of Recall@$K$ on $D_{full}$, which consists of the following steps:
@@
  • Compute the $\widehat{\text{Recall@}K}$ for this user: $\widehat{\text{Recall@}K}=\frac{m}{n},$ where $n$ denotes the number of items with positive label in Sorted $D_{rand}$, and $m$ denotes the number of these items with higher $\hat{y}$ than the ($K$+1)-th item.
@@
(11) $E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall@}K}\right]=\text{Recall}@K,$
@@
This signifies that the estimates derived from URE are unbiased.
@@
Since Recall is only related to positive samples, we can only consider positive samples. Let’s assume that in $D_{full}$, there are a total of $N$ positive samples, out of which $M$ samples are ranked higher than the sample ranked at $K+1$ in Sorted ItemID in Figure 2. Assuming that we have randomly sampled $\overline{N}$ samples, which have $n$ positive samples and the rest are negatives. Since the sampling process is random, the probability of the number $m$ of positive samples that rank higher than the sample ranked at $K+1$ in Sorted ItemID can be calculated as $(C^{m}_{M}C^{n-m}_{N-M}/C^{n}_{N})$, thus we have:
Figure 2. Illustration of the proposed URE scheme on unbiasedly estimating Recall@$K$ on $D_{full}$.
@@
(12) $E_{\mathcal{S}_{\overline{N}}}\left[\widehat{\text{Recall}@K}\right]$ $=E_{n^{\prime}}\left[E_{\mathcal{S}_{\overline{N}}}[m/n|n^{\prime}]\right]$
(13) $=E_{n^{\prime}}\left[\sum_{{m=1}}^{min(M,n^{\prime})}\frac{mn^{\prime}M}{mn^{\prime}N}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$
(14) $=E_{n^{\prime}}\left[\frac{M}{N}\sum_{{m=1}}^{min(M,n^{\prime})}\frac{C^{m-1}_{M-1}C^{n^{\prime}-1-(m-1)}_{N-M}}{C^{n^{\prime}-1}_{N-1}}\right]$
(15) $=M/N=\text{Recall}@K.$
Figure 3. (a) The correlation coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ as well as the conventional Recall@$\overline{K}$. (b) The correlation coefficients between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$.
5. Experiments
Unbiasedness Analysis. In this subsection, we test the AutoDebias model using the KuaiRec dataset to verify that the $\widehat{\text{Recall@}K}$ in the URE scheme is an unbiased estimation of Recall@$K$ on $D_{full}$.
In Figure 3a, we calculate the correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$, as well as Recall@$\overline{K}$ to demonstrate the ranking consistency. As shown in Figure 3a, when $K$ is small (e.g. $K=30$), both Recall@$\overline{1}$ and Recall@$\overline{5}$ exhibits poor correlations with Recall@$K$. In contrast, the correlation coefficient between the estimated $\widehat{\text{Recall@}K}$ and Recall@$K$ is higher than 0.9. This means that the model we choose based on $\widehat{\text{Recall@}K}$ is more likely to perform better on Recall@$K$ on $D_{full}$, which further demonstrates the effectiveness of the URE scheme.
In Figure 3b, we report the correlation coefficient between Recall@$K$ and $\widehat{\text{Recall@}K}$ for different values of $K$. As we desired, only when K is the same, Recall@$K$ and $\widehat{\text{Recall@}K}$ have the highest correlation coefficients. That is, URE allows us to select the best debiased model based on the Recall@$K$ with a desired $K$.
Based on our discussion, we show that our URE scheme unbiasedly estimates Recall@$K$ for different values of $K$ and ensure an accurate debiased model selection.
Table 1. The evaluation results by URE. Notably, the Recall@$5$ is not available in Yahoo!R3, as it lacks a fully-exposed dataset. The results are statistically significant with $p<$ 0.01.
Model KuaiRec Yahoo!R3
Traditional scheme Fully-Exposed Dataset URE scheme Traditional scheme URE scheme
Recall@$\overline{5}$ Recall@$5$ $\widehat{\text{Recall@5}}$ Recall@$\overline{5}$ $\widehat{\text{Recall@5}}$
MF 0.4859 0.0257 0.0267 0.7085 0.0131
IPS 0.4889 0.0266 0.0273 0.6572 0.0095
DR 0.4849 0.0262 0.0267 0.7195 0.0075
AutoDebias 0.5181 0.0317 0.0323 0.8085 0.0550
DNS 0.5288 0.0237 0.0245 0.8330 0.0490
Re-evaluation Results. In this subsection, we re-evaluate some classical debiasing methods under backbone MF (Koren et al., 2009), including IPS (Schnabel et al., 2016), DR (Wang et al., 2019), AutoDebias (Chen et al., 2021), DNS (Shi et al., 2023) on two widely-used datasets (cf. Table 2) in the debiasing research domain, and report the results in Table 1. We have the following key findings: (1) Based on the results on KuaiRec, we observe that the evaluation result $\widehat{\text{Recall@5}}$ of the URE scheme aligns closely with the true Recall@5 on $D_{full}$. This observation suggests a strong unbiasedness in the URE scheme. In contrast, the evaluation results of the traditional evaluation scheme exhibit a significant discrepancy from the true Recall@5, and it fails to ensure the preservation of order in the results. (2) When using $\widehat{\text{Recall@5}}$ to verify the debiasing performance, some debiasing methods perform worse than the orignal MF, which encourages the development of further debiasing techniques.
Table 2. The statistics of datasets.
Dataset ##\##Users ##\##Items Biased train test
Yahoo!R3 15,400 1,000 311,704 48,600 ($D_{rand}$)
KuaiRec 1411 3327 1,934,404 2,742,166 ($D_{full}$)
6. Conclusion
In this paper, we have unveiled the inconsistency between traditional evaluation scheme on the $D_{rand}$ and the gold standard evaluation on the $D_{full}$, which results in an inaccurate assessment of debiasing methods. To tackle this issue, we have introduced the URE scheme to unbiasedly estimate true Recall performance. Our proposal has been substantiated through both theoretical and empirical analyses, affirming its effectiveness.
Acknowledgements.
This work is supported by the National Key Research and Development Program of China (2022YFB3104701) and the National Natural Science Foundation of China (62272437).